OpenStudy (anonymous):

You are given functions f and g such that f(n)=O(g(n)). Is f(n)∗log2(f(n)c)=O(g(n)∗log2(g(n))) ? (Here c is some constant >0. You can assume that f and g are always bigger than 1. a) true b)false c)depends on c d) depends on f and g

OpenStudy (anonymous):

I think could go this way, but not sure: f(n)=O(g(n)) means |f(n)|<= M |g(n)| or |f(n)|/|g(n)| <= M so you have to prove, given that f(n)=O(g(n)), that |f(n)∗log2(f(n)c)|/|g(n)∗log2(g(n))| <= N |f(n)∗log2(f(n)c)|/|g(n)∗log2(g(n))| <= M|log2f(n)c)|/|log2(g(n))| = =M log2|(f(n)c -g(n))| <= M log2 (|M|g(n)|c - g(n)|). So i think it depends on c.